OpenStudy (anonymous):

3x+5=0 x= can you show me the steps with detail thxs

OpenStudy (anonymous):

3x+5=0 -5. -5 3x=-5 /3. /3 X=-3/5

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

yah\

OpenStudy (anonymous):

uh.... i think that is wrong?

OpenStudy (anonymous):

-5/3

OpenStudy (anonymous):

yes OMG, -5/3 is the correct answer

OpenStudy (anonymous):

Oh yeah, that part was flipped on my ey. Fr sure

OpenStudy (anonymous):

thats the answer i just wanted to know the steps

OpenStudy (anonymous):

thxs